Circular No. 7371

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000K IN MCG +9-19-191 Independent discoveries of a supernova in MCG +9-19-191 have been

reported as a result of unfiltered CCD images taken by Paolo Sala, Munich, on Feb. 28.07 UT (mag ∼ 16.6; instrument not specified) and by T. Puckett, Mountain Town, GA, and D. George, Ottawa, ON, on Feb. 28.36 (mag 18.1) with the Puckett Observatory 0.35-m automated supernova patrol telescope (limiting mag 19.5). Sala provides the position of SN 2000K as α = 11^{h}47^{m}49^{s}.76, δ = +56^{o}01′00′′.1 (equinox 2000.0), which is ∼ 15′′ east-southeast of the host galaxy, noting that this location is ∼ 200′′ northnortheast of NGC 3888. Puckett gives the position end figures for the supernova as 49^{s}.99, 01′′.4, or 15′′.2 east and 5′′.6 south of the center of MCG +9-19-191. Sala notes that the new star is absent on a comparison image of the galaxy obtained on 1998 Apr. 22. Puckett reports that SN 2000K was also present on an unfiltered CCD frame taken on 2000 Feb. 29.08, but it was not present on a frame taken on Feb. 9.28, and it does not appear on Palomar Sky Survey images taken on 1993 Apr. 25 (limiting mag ∼ 21.0) or 1950 Mar. 21 (limiting mag ∼ 20.0).

S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that images of SN 2000K, taken on Feb. 29.4 UT by P. Barmby with the F. L. Whipple Observatory (FLWO) 1.2-m telescope (+ 4shooter), yield the following precise position for SN 2000K: α = 11^{h}47^{m}50^{s}.08, δ = +56^{o}01′01′′.3 (equinox 2000.0), which is 15′′.5 east and 4′′.8 south of the host galaxy. They add: “Concurrent spectra of SN 2000K, taken by P. Berlind with the FLWO 1.5-m telescope (+ FAST spectrograph), show it to be a type-Ia supernova after maximum light. The age of the supernova, based on its spectral features (Riess et al. 1997, A.J. 114,

722) is 13 ± 3 days past maximum light. If we adopt a Si II (rest 635.5 nm) expansion velocity of 10 000 km/s, typical of type-Ia supernovae at this epoch (Jha et al. 1999, Ap.J. Supp. 125, 73), the observed Si II feature implies a redshift of 16 000 km/s for the host galaxy.”

SUPERNOVA 1999cm IN UGC 9766 Corrigendum. On IAUC 7190, line 5, for 1′′.0 east read 7′′.5 east

2000 February 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green